# Order-Optimal Instance-Dependent Bounds for Offline Reinforcement Learning with Preference Feedback

### v1

@@ Appendix C The existence of $\hat{i}_{k}$ in RL-LOW and DP-RL-LOW
Corollary C.2.
For any consistent instance $v$ and using the estimate of the best action under each state $k$ as prescribed by RL-LOW-DP, we have $\hat{i}_{k}$ exists.
Appendix D Proof of Lower Bound
@@
where ${\rm Sig}(\cdot)$ represents the Sigmoid function. By the fact that $-2R_{\rm max}\leq\langle\phi(k,i)-\phi(k,j),\theta\rangle\leq 2R_{\rm max}$ and $-2R_{\rm max}\leq\langle\phi(k,i)-\phi(k,j),\theta^{\prime}\rangle\leq 2R_{\rm max}$ and that $\frac{\mathrm{d}{\rm Sig}(x)}{\mathrm{d}x}=\frac{\exp(-x)}{(\exp(-x)+1)^{2}}$, we further get that
@@
Then, by Corollary D.1, we get that Eqn. (36) and Eqn. (37) imply that
@@
Then, recall that we denote $\tilde{D}(\cdot,\cdot)$ as the approximation of KL divergence between instance $v$ and $v^{\prime}$ that share the same parameter except $\theta$, i.e., for any $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$ and $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta^{\prime})$,
@@
We are ready to prove Lemma 4.2 that reveals a useful property for ${\rm Alt}(\cdot)$ and $\tilde{D}(\cdot)$
@@
Let $\mathcal{G}$ be an arbitrary orthonormal basis of ${\rm Span}\{\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime}):(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}^{2}\text{ and }N_{k^{\prime},i^{\prime},j^{\prime}}>0\}$. Also let $[\mathbf{w}]_{\mathcal{G}}$ denote the column vector that represents $\mathbf{w}$ under basis $\mathcal{G}$ (Meyer,, 2000, Chapter 4). Define the matrix
@@
$\begin{cases}\lambda=\frac{2\eta}{\lVert\mathbb{[z]_{\mathcal{G}}}\rVert^{2}_{V^{-1}}}\\ \mathbb{x}=-\frac{\eta}{\lVert[\mathbb{z}]_{\mathcal{G}}\rVert^{2}_{V^{-1}}}V^{-1}\mathbb{z}+\mathbb{g},\\ \end{cases}$
@@
$\begin{cases}\langle\phi(k,i^{*}_{k}(v))-\phi(k,i),\theta-\theta^{\prime}\rangle=2\Delta_{k,i}(v)\qquad\mbox{for}\ (k,i)=(\bar{k}(v),\bar{i}(v))\\ \lvert\langle\phi(k,i^{*}_{k}(v))-\phi(k,i),\theta-\theta^{\prime}\rangle\rvert\leq\frac{1}{2}\Delta_{k,i}(v)\qquad\forall i\neq i^{*}_{k}(v)\mbox{ and}(k,i)\neq(\bar{k}(v),\bar{i}(v))\\ \end{cases}$ (42)
@@
$\begin{cases}i^{*}_{k}(v)=i^{*}_{k}(v^{\prime})\quad\forall k\neq\bar{k}(v)\\ \bar{k}(v)=\bar{k}(v^{\prime})\\ i^{*}_{\bar{k}(v^{\prime})}(v^{\prime})=\bar{i}(v)\\ \end{cases}$ (43)
@@
Fix any $v$ and $v^{\prime}$ as in Lemma D.2. By the definition of $v^{\prime}$, we can immediately obtain the first inequality of Eqn. (42), i.e.,
@@
<$\stackrel{{\scriptstyle(a)}}{{<}}16\tilde{D}(v,v^{\prime})$
<$\stackrel{{\scriptstyle(b)}}{{<}}\frac{16\Delta_{\bar{k}(v),\bar{i}(v)}(v)^{2}}{\lVert[\phi(\bar{k}(v),\bar{i}(v))-\phi(\bar{k}(v),i^{*}_{\bar{k}(v)})]_{\mathcal{G}}\rVert^{2}_{V^{-1}}}$
<$\stackrel{{\scriptstyle(c)}}{{<}}\frac{64}{\gamma_{\bar{k}(v),\bar{i}(v)}(v)}$ (47)
where (a) follows from Eqn. (45) and (b) follows from Eqn. (46), and (c) follows from Lemma C.1.
@@
By the fact that $v^{\prime\prime}\in{\rm Alt}(v,\phi(\tilde{k},\tilde{i})-\phi(\tilde{k},i^{*}_{\tilde{k}}))$ as well as Eqn. (24) and Eqn. (25), we further get that
@@
$\begin{cases}i^{*}_{k}(v)=i^{*}_{k}(v^{\prime})\quad\forall k\neq\bar{k}(v)\\ \bar{k}(v)=\bar{k}(v^{\prime})\\ i^{*}_{\bar{k}(v^{\prime})}(v^{\prime})=\bar{i}(v)\\ \end{cases}$
This completes the proof of Lemma D.2. ∎
Lemma D.3.
@@
Fix any $v$ and $v^{\prime}$ as in Lemma D.2. Note that from Lemma D.2, we get that instance $v$ and instance $v^{\prime}$ share the same optimal action for the state $k\neq\bar{k}$ and the optimal action of state $\bar{k}(v)$ is $\bar{i}(v)$ under the instance $v^{\prime}$, i.e., $i^{*}_{\bar{k}(v^{\prime})}(v^{\prime})=\bar{i}(v)$ and $i^{*}_{k}(v)=i^{*}_{k}(v^{\prime})$ for all $k\neq\bar{k}(v)$. Then, by definition, the hardness of instance $v^{\prime}$ is
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\max\left\{\max_{k\in\mathcal{S},i\in\mathcal{A}:i\neq i^{*}_{k}(v^{\prime}),k\neq\bar{k}(v)}\ \frac{\gamma_{k,i}(v)}{\Delta_{k,i}^{2}(v^{\prime})},\max_{i\in\mathcal{A}:i\neq\bar{i}(v),k=\bar{k}(v)}\frac{\gamma_{k,i}(v^{\prime})}{\Delta_{k,i}^{2}(v^{\prime})}\right\}$
@@
Then, by the fact that $\lvert\langle\phi(k,i^{*}_{k}(v))-\phi(k,i),\theta-\theta^{\prime}\rangle\rvert\leq\frac{1}{2}\Delta_{k,i}(v)$ for all $i\neq i^{*}_{k}(v)\text{ and }(k,i)\neq(\bar{k}(v),\bar{i}(v))$ from Lemma D.2, we get that
@@
Further, by the fact that $\langle\phi(k,i^{*}_{k}(v))-\phi(k,i),\theta-\theta^{\prime}\rangle=2\Delta_{k,i}(v)$ for $(k,i)=(\bar{k}(v),\bar{i}(v))$ Lemma D.2, we get that
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\max\left\{H(v),\max_{i\in\mathcal{A}:i\neq\bar{i}(v),i\neq i^{*}_{\bar{k}(v)}(v)}\frac{\lVert[\phi(\bar{k}(v),\bar{i}(v))-\phi(\bar{k}(v),i)]_{\mathcal{G}}\rVert^{2}_{V^{-1}}}{\Delta_{\bar{k}(v),i}^{2}(v^{\prime})}\right\}$ (50)
where (a) follows from Lemma C.1, $V$ and $\mathcal{G}$ are as defined in Lemma 4.2. Similarly, by the fact that $\lvert\langle\phi(k,i^{*}_{k}(v))-\phi(k,i),\theta-\theta^{\prime}\rangle\rvert\leq\frac{1}{2}\Delta_{k,i}(v)\ \ \forall i\neq i^{*}_{k}(v)\text{ and }(k,i)\neq(\bar{k}(v),\bar{i}(v))$, we obtain for $i\in\mathcal{A}$ with $i\neq\bar{i}(v),i\neq i^{*}_{\bar{k}(v)}(v)$
@@
Finally, combining Eqn. (D), Eqn. (50) and Eqn. (51), we complete the proof of Lemma D.3. ∎
@@
By Lemma D.3, we get that $H(v^{\prime})\leq H(v)\leq 8H(v^{\prime})$. By Lemma 4.2, we get that
@@
where $\mathcal{G}$ and $V$ are as defined in Lemma 4.2. Further, by Lemma 4.1 and Lemma D.2, we get that for any $n>0$
@@
By the definitions of $v$ and $v^{\prime}$, we can obtain that $\bar{\Delta}>0$. Then, we get that for any policy $\pi$
@@
where $D_{\rm TV}(\cdot,\cdot)$ denotes the total variance distance, and (a) follows from Bretagnolle–Huber inequality (Tsybakov,, 2009, Lemma 2.6) and (c) follows from Eqn. (52). Finally, for all sufficiently large $n$, we have
@@
Appendix E Proof of Upper bound
Lemma E.1.
@@ Appendix E Proof of Upper bound
Proof of Lemma E.1.
For simplicity of notation, we let $x_{0}\coloneqq\mathbb{E}(X_{n})$, which implies $x_{0}=p\in[1-\beta,\beta]$. Then, by the smoothness of $f(\cdot)$ on $[1-\beta,\beta]$, we obtain the equivalent expression of $f(\cdot)$ from the Taylor expansion of $f(\cdot)$ on $[1-\beta,\beta]$,
$f(x)=\begin{cases}f(x_{0})+f^{\prime}(x_{0})\cdot(x-x_{0})+\frac{1}{2}f^{\prime\prime}(\xi_{x})\cdot(x-x_{0})^{2}\quad\text{ if }x\in[1-\beta,\beta]\\ f(x_{0})+f^{\prime}(x_{0})\cdot(\beta-x_{0})+\frac{1}{2}f^{\prime\prime}(\xi_{\beta})\cdot(\beta-x_{0})^{2}\quad\text{ if }x>\beta\\ f(x_{0})+f^{\prime}(x_{0})\cdot(1-\beta-x_{0})+\frac{1}{2}f^{\prime\prime}(\xi_{\beta})\cdot(1-\beta-x_{0})^{2}\quad\text{ if }x<1-\beta,\\ \end{cases}$ (55)
@@
where ${\rm Var}(\cdot)$ represents the variance of the random variable $\cdot$ and $[x]_{+}=\max\{x,0\}$.
@@
Then, by Lemma A.5, we get that
@@
Similarly, from Lemma A.5, we further get that
@@
Note that by definition, we have $\|f(X_{n})-f(p)\|_{\rm vp}^{2}=\|f(X_{n})\|_{\rm vp}^{2}$, which completes the proof of Lemma E.2. ∎
@@
Fix any consistent instance $v$ and dataset size $n$ under RL-LOW. By Lemma E.2 and Lemma A.3, we get that $\hat{r}_{k,i^{*}_{k},i}$ is subGaussian with variance proxy as
@@
In addition, by Lemma E.1, we get that for any $(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{A}$ with $N_{k^{\prime},i^{\prime},j^{\prime}}>0$,
@@
$R_{n}$ $\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\mathbbm{1}_{\{\hat{r}_{k,i^{*}_{k},i}\leq 0\}}\cdot\rho_{k}\Delta_{k,i}$ (66)
$\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\bigg{[}\bigg{(}\mathbbm{1}_{\{\hat{r}_{k,i^{*}_{k},i}\leq\mathbb{E}_{v}^{\rm RL\mbox{-}LOW}\left[\hat{r}_{k,i,i^{*}_{k}}\right]-\Delta_{k,i}/2\}}$
$\hskip 85.35826pt\lor\mathbbm{1}_{\{\lvert\mathbb{E}_{v}^{\rm RL\mbox{-}LOW}\left[\hat{r}_{k,i^{*}_{k},i}\right]-\Delta_{k,i}\rvert\geq\Delta_{k,i}/2\}}\bigg{)}\cdot\rho_{k}\Delta_{k,i}\bigg{]}$ (67)
By Lemma A.4 and Eqn. (62), we get that for any $(k,i)\in\mathcal{S}\times\mathcal{A}$ with $i\neq i^{*}_{k}$,
@@
$R_{n}$ $\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\mathbbm{1}_{\{\hat{r}_{k,i^{*}_{k},i}\leq 0\}}\cdot\rho_{k}\Delta_{k,i}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\rho_{k}\bigg{[}\mathbbm{1}_{\{\Delta_{k,i}<\frac{6\tilde{\gamma}_{k,i}}{(1-\beta)^{4}\sqrt{n}}\}}\cdot\frac{6\tilde{\gamma}_{k,i}}{(1-\beta)^{4}\sqrt{n}}$
$\hskip 85.35826pt+\mathbbm{1}_{\{\lvert\mathbb{E}_{v}^{\rm RL\mbox{-}LOW}\left[\hat{r}_{k,i,i^{*}_{k}}\right]-r_{k,i,i^{*}_{k}}\rvert\leq\Delta_{k,i}/2\}}\cdot\Delta_{k,i}\bigg{]}$

### v2

@@ Appendix D Useful facts
Corollary F.2.
For any consistent instance $v$ and using estimate of the best action under each state $k$ as prescribed by DP-RL-LOW, we have $\lvert\tilde{\mathcal{I}}_{k}\rvert\geq 1$ and
$\operatorname*{argmax}_{i\in\mathcal{A}}\tilde{r}_{k,i,j_{1}}=\operatorname*{argmax}_{i\in\mathcal{A}}\tilde{r}_{k,i,j_{2}}=\tilde{\mathcal{I}}_{k}\quad\mbox{for any}\quad j_{1},j_{2}\in\mathcal{A}.$ (35)
F.2 Computational Complexity
computational complexity of RL-LOW and DP-RL-LOW
With Proposition 3.2, it is not necessary to compute all values of $\hat{r}_{k,i,j}$ for each $(k,i,j)\in\mathcal{S}\times\mathcal{A}^{2}$. Instead, we only need to compute $\hat{r}_{k,i,j^{\dagger}}$ for each $(k,i)\in\mathcal{S}\times\mathcal{A}$ where $j^{\dagger}\in\mathcal{A}$ is (arbitrarily) fixed (e.g., $j^{\dagger}=1$), and let $\hat{i}_{k}\in\arg\max_{i\in\mathcal{A}}\hat{r}_{k,i,j^{\dagger}}$.
Note that by Lemma F.1, we have
$\hat{r}_{k,i,j^{\dagger}}=\sum_{(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}^{2}}N_{{k^{\prime}},{i^{\prime}},{j^{\prime}}}[\phi({k^{\prime}},{i^{\prime}})-\phi({k^{\prime}},{j^{\prime}})]_{\mathcal{G}}^{\top}V^{-1}[\phi(k,i)-\phi(k,j^{\dagger})]_{\mathcal{G}}\log\Big(\frac{B^{\rm CLP}_{k^{\prime},i^{\prime},j^{\prime}}}{1-B^{\rm CLP}_{k^{\prime},i^{\prime},j^{\prime}}}\Big).$
Then, if we pre-calculate a global vector of $\sum_{(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}^{2}}N_{{k^{\prime}},{i^{\prime}},{j^{\prime}}}[\phi({k^{\prime}},{i^{\prime}})-\phi({k^{\prime}},{j^{\prime}})]_{\mathcal{G}}^{\top}V^{-1}\log\Big(\frac{B^{\rm CLP}_{k^{\prime},i^{\prime},j^{\prime}}}{1-B^{\rm CLP}_{k^{\prime},i^{\prime},j^{\prime}}}\Big)$, we can compute each $\hat{r}_{k,i,j^{\dagger}}$ for $(k,i)\in\mathcal{S}\times\mathcal{A}$ in $\mathcal{O}(d)$ time complexity. Similarly, $\tilde{r}_{k,i,j^{\dagger}}$ in DP-RL-LOW can be computed through an analogous procedure. Hence, the overall computational complexity of RL-LOW and DP-RL-LOW are $\mathcal{O}(SAd+nd^{2}+d^{3})$, where the term “SAdSAd” corresponds to the natural process of scanning the feature vectors for all state-action pairs, and the terms $nd^{2}+d^{3}$” corresponds to compute $V^{-1}$ in above.
The computational complexity’s dependence on “$SAd$ ” is inevitable in our current framework, as the generation of the output $\hat{i}_{k}$ is required for all $k\in\mathcal{S}$. Nonetheless, in an alternative setting in which the set of best actions to be estimated $\{\hat{i}_{k}\in\mathcal{A}\}_{k\in\mathcal{S}}$ is replaced by a parametric function $\{\hat{i}(k;\vartheta)\in\mathcal{A}\}$ where a parameter $\vartheta$ is to be estimated. In this setting, the previously global vector can be utilized to represent $\vartheta$. Then, in this setting, the overall computational complexity becomes $O(nd^{2}+d^{3})$, which is required to compute the global vector.
computational complexity of RL-LOW-MDP
In RL-LOW-MDP, after each $\hat{r}_{k,i,j^{\dagger}}$ for $(k,i)\in\mathcal{S}\times\mathcal{A}$ is computed, the rest process is the standard RL problem under the condition that the transition and reward functions are known. Hence, the overall computational complexity of RL-LOW-MDP is $\mathcal{O}(SAd+nd^{2}+d^{3}+g(S,A))$, where the term $g(S,A)$ corresponds to the above standard problem that can be solved by asynchronous dynamic programming or linear programming (Sutton and Barto 2018). Particularly $g(S,A)=O(\kappa SA)$ by using asynchronous dynamic programming, and $\kappa$ is a hyperparameter that represents the average number of iteration steps, which controls the solution precision.
Appendix G Proof of Lower Bound
@@
where ${\rm Sig}(\cdot)$ represents the Sigmoid function. By the fact that $2R_{\rm max}\leq\langle\phi(k,i)-\phi(k,j),\theta\rangle\leq 2R_{\rm max}$ and $2R_{\rm max}\leq\langle\phi(k,i)-\phi(k,j),\theta^{\prime}\rangle\leq 2R_{\rm max}$ and that $\frac{\mathrm{d}{\rm Sig}(x)}{\mathrm{d}x}=\frac{\exp(-x)}{(\exp(-x)+1)^{2}}$, we further get that
@@
Then, by Corollary G.1, we get that Eqn. (39) and Eqn. (40) imply that
@@
[…]
@@
[…]
@@
Let $\mathcal{G}$ be an arbitrary orthonormal basis of ${\rm Span}\{\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime}):(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}^{2}\text{ and }N_{k^{\prime},i^{\prime},j^{\prime}}>0\}$. Also let $\mathbf{w}]_{\mathcal{G}}$ denote the column vector that represents $\mathbf{w}$ under basis $\mathcal{G}$ (Meyer 2000, Chapter 4). Define the matrix
@@
[…]
@@
[…]
@@
[…]
@@
Fix any $v$ and $v^{\prime}$ as in Lemma G.2. By the definition of $v^{\prime}$, we can immediately obtain the first inequality of Eqn. (45), i.e.,
@@
$\stackrel{{\scriptstyle(a)}}{{<}}16\tilde{D}(v,v^{\prime})$
$\stackrel{{\scriptstyle(b)}}{{<}}\frac{16\Delta_{\bar{k}(v),\bar{i}(v)}(v)^{2}}{\lVert[\phi(\bar{k}(v),\bar{i}(v))-\phi(\bar{k}(v),i^{*}_{\bar{k}(v)})]_{\mathcal{G}}\rVert^{2}_{V^{-1}}}$
$\stackrel{{\scriptstyle(c)}}{{<}}\frac{64}{\gamma_{\bar{k}(v),\bar{i}(v)}(v)}$ (50)
where (a) follows from Eqn. (48) and (b) follows from Eqn. (49), and (c) follows from Lemma F.1.
@@
By the fact that $)v^{\prime\prime}\in{\rm Alt}(v,\phi(\tilde{k},\tilde{i})-\phi(\tilde{k},i^{*}_{\tilde{k}}$)) as well as Eqn. (25) and Eqn. (26), we further get that
@@
[…]
This completes the proof of Lemma G.2. ∎
Lemma G.3.
@@
Fix any $v$ and $v^{\prime}$ as in Lemma G.2. Note that from Lemma G.2, we get that instance $v$ and instance $v^{\prime}$ share the same optimal action for the state $k\neq\bar{k}$ and the optimal action of state $\bar{k}(v)$ is $\bar{i}(v)$ under the instance $v^{\prime}$, i.e., $i^{*}_{\bar{k}(v^{\prime})}(v^{\prime})=\bar{i}(v)$ and $i^{*}_{k}(v)=i^{*}_{k}(v^{\prime})$ for all $k\neq\bar{k}(v)$. Then, by definition, the hardness of instance $v^{\prime}$ is
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\max\left\{\max_{k\in\mathcal{S},i\in\mathcal{A}:i\neq i^{*}_{k}(v^{\prime}),k\neq\bar{k}(v)}\ \frac{\gamma_{k,i}(v)}{\Delta_{k,i}^{2}(v^{\prime})},\max_{i\in\mathcal{A}:i\neq\bar{i}(v),k=\bar{k}(v)}\frac{\gamma_{k,i}(v^{\prime})}{\Delta_{k,i}^{2}(v^{\prime})}\right\}$
@@
Then, by the fact that $\lvert\langle\phi(k,i^{*}_{k}(v))-\phi(k,i),\theta-\theta^{\prime}\rangle\rvert\leq\frac{1}{2}\Delta_{k,i}(v)$ for all $i\neq i^{*}_{k}(v)\text{ and }(k,i)\neq(\bar{k}(v),\bar{i}(v))$ from Lemma G.2, we get that
@@
Further, by the fact that $\langle\phi(k,i^{*}_{k}(v))-\phi(k,i),\theta-\theta^{\prime}\rangle=2\Delta_{k,i}(v)$ for $(k,i)=(\bar{k}(v),\bar{i}(v))$ Lemma G.2, we get that
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\max\left\{H(v),\max_{i\in\mathcal{A}:i\neq\bar{i}(v),i\neq i^{*}_{\bar{k}(v)}(v)}\frac{\lVert[\phi(\bar{k}(v),\bar{i}(v))-\phi(\bar{k}(v),i)]_{\mathcal{G}}\rVert^{2}_{V^{-1}}}{\Delta_{\bar{k}(v),i}^{2}(v^{\prime})}\right\}$ (53)
where (a) follows from Lemma F.1, $V$ and $\mathcal{G}$ are as defined in Lemma 4.2. Similarly, by the fact that $\lvert\langle\phi(k,i^{*}_{k}(v))-\phi(k,i),\theta-\theta^{\prime}\rangle\rvert\leq\frac{1}{2}\Delta_{k,i}(v)\ \ \forall i\neq i^{*}_{k}(v)\text{ and }(k,i)\neq(\bar{k}(v),\bar{i}(v))$, we obtain for $i\in\mathcal{A}$ with $i\neq\bar{i}(v),i\neq i^{*}_{\bar{k}(v)}(v)$
@@
Finally, combining Eqn. (G), Eqn. (53) and Eqn. (54), we complete the proof of Lemma G.3. ∎
@@
By Lemma G.3, we get that $H(v^{\prime})\leq H(v)\leq 8H(v^{\prime})$. By Lemma 4.2, we get that
@@
where $\mathcal{G}$ and $V$ are as defined in Lemma 4.2. Further, by Lemma 4.1 and Lemma G.2, we get that for any $n>0$
@@
By the definitions of $v$ and $v^{\prime}$, we can obtain that $\bar{\Delta}>0$. Then, we get that for any algorithm $\Pi$
@@
where $D_{\rm TV}(\cdot,\cdot)$ denotes the total variance distance, and (a) follows from Bretagnolle–Huber inequality (Tsybakov 2009, Lemma 2.6) and (c) follows from Eqn. (55). Finally, for all sufficiently large $n$, we have
@@
Appendix H Proof of Upper bounds
Lemma H.1.
@@ Appendix H Proof of Upper bounds
Proof of Lemma H.1.
For simplicity of notation, we let $x_{0}\coloneqq\mathbb{E}(X_{n})$, which implies $x_{0}=p\in[1-\beta,\beta]$. Then, by the smoothness of $f(\cdot)$ on $1-\beta,\beta]$, we obtain the equivalent expression of $f(\cdot)$ from the Taylor expansion of $f(\cdot)$ on $1-\beta,\beta]$,
[…]
@@
where ${\rm Var}(\cdot)$ represents the variance of the random variable $\cdot$ and $x]_{+}=\max\{x,0\}$.
@@
Then, by Lemma D.5, we get that
@@
Similarly, from Lemma D.5, we further get that
@@
Note that by definition, we have $\|f(X_{n})-f(p)\|_{\rm vp}^{2}=\|f(X_{n})\|_{\rm vp}^{2}$, which completes the proof of Lemma H.2. ∎
@@
Fix any consistent instance $v$ under RL-LOW. By Lemma H.2 and Lemma D.3, we get that $\hat{r}_{k,i^{*}_{k},i}$ is subGaussian with variance proxy as
@@
In addition, by Lemma H.1, we get that for any $(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{A}$ with $N_{k^{\prime},i^{\prime},j^{\prime}}>0$,
@@
$R_{n}$ $\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\mathds{1}_{\{\hat{r}_{k,i^{*}_{k},i}\leq 0\}}\cdot\rho_{k}\Delta_{k,i}$ (69)
$\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\bigg[\bigg(\mathds{1}_{\{\hat{r}_{k,i^{*}_{k},i}\leq\mathbb{E}_{v}^{\rm RL\mbox{-}LOW}\left[\hat{r}_{k,i,i^{*}_{k}}\right]-\Delta_{k,i}/2\}}$
$\hskip 85.35826pt\lor\mathds{1}_{\{\lvert\mathbb{E}_{v}^{\rm RL\mbox{-}LOW}\left[\hat{r}_{k,i^{*}_{k},i}\right]-\Delta_{k,i}\rvert\geq\Delta_{k,i}/2\}}\bigg)\cdot\rho_{k}\Delta_{k,i}\bigg]$ (70)
By Lemma D.4 and Eqn. (65), we get that for any $(k,i)\in\mathcal{S}\times\mathcal{A}$ with $i\neq i^{*}_{k}$,
@@
$R_{n}$ $\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\mathds{1}_{\{\hat{r}_{k,i^{*}_{k},i}\leq 0\}}\cdot\rho_{k}\Delta_{k,i}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\rho_{k}\bigg[\mathds{1}_{\{\Delta_{k,i}<\frac{6\tilde{\gamma}_{k,i}}{(1-\beta)^{4}\sqrt{n}}\}}\cdot\frac{6\tilde{\gamma}_{k,i}}{(1-\beta)^{4}\sqrt{n}}$
$\hskip 85.35826pt+\mathds{1}_{\{\lvert\mathbb{E}_{v}^{\rm RL\mbox{-}LOW}\left[\hat{r}_{k,i,i^{*}_{k}}\right]-r_{k,i,i^{*}_{k}}\rvert\leq\Delta_{k,i}/2\}}\cdot\Delta_{k,i}\bigg]$